complete the table below by writing the correct symbols and formula in each parameter and statistic.

						       Parameter	    | 	   Statistic

Different Measures | Symbol | Formula | Symbol | Formula

  • MEAN | | | |
  • VARIANCE | | | |
  • STANDARD DEVIATION | | | |
1
Expert's answer
2021-06-03T15:44:52-0400

Table:





1.) Mean :

(a) Population mean:


Symbol : "\mu"

Formula : "\mu=\dfrac{\sum_{i=1}^N\ x_i}{N}"


(b) Sample mean:


Symbol: "\bar x"

Formula: "\bar x =\dfrac{\sum_{i=1}^n x_i}{n}"




2.) Variance:


(a) Population Variance


Symbol: "\sigma^2"

Formula: "\sigma^2=\dfrac{\sum_{i=1}^N (x_i-\mu)^2}{N}"


(b) Sample Variance


Symbol: "s^2"

Formula:"s^2=\dfrac{\sum_{i=1}^n (x_i-\bar x)^2}{n}"



(c) Standard Deviation


(a) Population Standard Deviation:


Symbol: "\sigma"

Formula: "\sigma=\sqrt{\dfrac{\sum_{i=1}^N (x_i-\mu)^2}{N}}"



(b) Sample Standard Deviation;


Symbol: "s"


Formula: "s=\sqrt{\dfrac{\sum_{i=1}^n (x_i-\bar x)^2}{n}}"